let

f(x) =1 + 2đ‘„, đ‘„ ≀ 0

=3đ‘„ − 2,0 < đ‘„ ≀ 1

= 2đ‘„ àŹ¶ − 1, đ‘„ > 1

i) Check whether f is discontinuous. If yes, find where? ii) Give a rough sketch of the graph of f. 


1
Expert's answer
2021-12-30T15:59:19-0500
"f(x) = \begin{cases}
 1+2x, x\leq 0 \\
 3x-2, 0<x\leq 1 \\
 2x-1, x> 1 \\
\end{cases}"

i)


"\lim\limits_{x\to 0^-}f(x)=\lim\limits_{x\to 0^-}(1+2x)=1+2(0)=1""\lim\limits_{x\to 0^+}f(x)=\lim\limits_{x\to 0^+}(3x-2)=3(0)-2=-2""\lim\limits_{x\to 0^-}f(x)=1\not=-2=\lim\limits_{x\to 0^+}f(x)""\lim\limits_{x\to 0}f(x) \text{does not exist}"

The function "f(x)" has an jump discontinuity at "x=0."



"\lim\limits_{x\to 1^-}f(x)=\lim\limits_{x\to 1^-}(3x-2)=3(1)-2=1""\lim\limits_{x\to 1^+}f(x)=\lim\limits_{x\to 1^+}(2x-1)=2(1)-1=1""\lim\limits_{x\to 1^-}f(x)=1=\lim\limits_{x\to 1^+}f(x)=>\lim\limits_{x\to 1}f(x)=1""f(1)=3(1)-2=1=\lim\limits_{x\to 1}f(x)"

The function "f(x)" is continuous at "x=1."


The function "f(x)" is discontinuous at "x=0."

The function "f(x)" has an jump discontinuity at "x=0."

The function "f(x)" is continuous at "x=1."


ii)